### v1

# Electric-Field Driven Nuclear Dynamics of Liquids and Solids from a Multi-Valued Machine-Learned Dipolar Model

Abstract
The driving of vibrational motion by external electric fields is a topic of continued interest, due to the possibility of assessing new or metastable material phases with desirable properties. Here, we combine ab initio molecular dynamics within the electric-dipole approximation with machine-learning neural networks (NNs) to develop a general, efficient and accurate method to perform electric-field-driven nuclear dynamics for molecules, solids, and liquids. We train equivariant and autodifferentiable NNs for the interatomic potential and the dipole, modifying the prediction target to account for the multi-valued nature of the latter in periodic systems. We showcase the method by addressing property modifications induced by electric field interactions in a polar liquid and a polar solid from nanosecond-long molecular dynamics simulations with quantum-mechanical accuracy. For liquid water, we present a calculation of the dielectric function in the GHz to THz range and the electrofreezing transition, showing that nuclear quantum effects enhance this phenomenon. For the ferroelectric perovskite LiNbO3, we simulate the ferroelectric to paraelectric phase transition and the non-equilibrium dynamics of driven phonon modes related to the polarization switching mechanisms, showing that a full polarization switch is not achieved in the simulations.
I Introduction
The interaction of matter with static and dynamical external electric fields plays a fundamental role in understanding the behavior of atoms, molecules, and complex materials. This interaction is exploited in fields as diverse as semiconductor nanotechnology [1] and enzymatic reactions [2]. In particular, there is a rising interest in tuning static and low-frequency dielectric fields to drive reversible phase transitions of materials, with the goal of controlling mechanical properties, increasing the efficiency of ionic conductors and capacitors, or making better energy or memory storage devices [1, 2, 3]. The large spatial extent and, depending on the phenomenon, the time-scale of many nanoseconds involved in the nuclear dynamics that define the response of a material to the application of such electric fields has been a long-standing challenge for quantum-mechanical first-principles atomistic simulations. While the physics of the matter-field coupling is known and appropriate simulation techniques already exist within the context of quantum-mechanical theories [4, 5, 6], simplified and non-transferable models [3] have often been used for describing these quantum phenomena at larger system sizes and longer time scales in complex systems.
The goal of this work is to achieve a general framework that: (i) is generally applicable to different material classes (e.g. molecules, solids, liquids, and disordered systems) without relying on system-specific assumptions; (ii) keeps quantum-mechanical accuracy for electronic interactions and can include nuclear quantum effects; (iii) allows performing nanoseconds-long simulations including thousands of atoms in and out of equilibrium. It is clear that the rapidly evolving techniques of machine-learning (ML) in atomistic simulations can be applied to this problem. ML models have already been proven to succeed in the prediction of energies, forces and electronic-structure properties of materials [7, 8] with an accuracy comparable to that of the underlying quantum electronic-structure method, but at a small fraction of the cost and often with a strongly reduced scaling with system size. A fundamental quantity to describe the interaction of matter with electric fields is the dipole and its derivative with respect to nuclear displacements. Several works have already proposed ML-based models for these quantities [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 13]. While these approaches are appropriate and successful in the investigated examples, i.e., close to structural equilibria, they can become inaccurate when large displacements take place. As we will show in this work, an additional term arising from the multi-valued nature of the dipole in periodic systems sensitively influences the applicability of the ML model in these cases.
By explicitly incorporating the multi-valued nature of the dipole in our model, we achieve a framework that is applicable to neutral and polar materials at different thermodynamic state points and far from equilibrium. We showcase the capabilities of our model by addressing liquid water and the solid ferroelectric perovskite LiNbO3 [20]. We calculate nanoseconds of first-principles-quality molecular dynamics coupled to static and time-dependent electric fields for these polar materials. For water, we show simulations of the dielectric function [21] spanning the GHz to the THz range and of the electrofreezing transition [22, 23], finding an enhancement of this phenomenon due to nuclear quantum effects at larger field intensities. For LiNbO3, we show the calculation of the ferroelectric to paraelectric phase transition [24, 25, 26, 27] and of the full-dimensional non-equilibrium and non-linear driving of phonon-modes in this material. For this last example, incorporating the multi-valued nature of the dipole in the ML model is essential. It shows that full-dimensional simulations can better explain the experimental observation that an ultrafast polarization switch is only transiently and partially achieved [28, 29, 30].
In the following, we will summarize the theory of the model we propose, explain the ML model we develop, and discuss the rich phenomenology that these simulations can address together with the novel physical insights they can bring.
II Theory of Molecular Dynamics Driven by Electric Fields
When the applied electric field $\bm{\mathcal{E}}$ is small compared to the internal fields, one can apply the electric dipole approximation (EDA) within the linear-response regime (see Supplementary Section S1). In this case, the nuclear Hamiltonian can be expressed as [4, 6, 5]
[…]
where $T(\bm{p})$ is the kinetic energy, $\bm{p}$ the nuclear momenta, $\bm{R}$ the nuclear coordinates, $V(\bm{R})\equiv E[n_{0}]$ is the Born-Oppenheimer potential energy surface (PES) obtained from the ground-state electronic density $n_{0}$, and $\bm{\mu}(\bm{R})\equiv\bm{\mu}[n_{0}]$ is the system dipole.
The forces on the nuclei depend on the external field $\bm{\mathcal{E}}$ through
[…]
The first term on the right-hand side is the usual Born-Oppenheimer (BO) force. The second term can be expressed in terms of the atomic-polar tensors [17, 31], also called Born effective charges (BEC) $\bm{Z}^{*}$,
$\bm{Z}^{*}=\,\frac{1}{e}\frac{\partial\bm{\mu}}{\partial\bm{R}}=\frac{1}{e}\frac{\partial\bm{F}}{\partial\bm{\mathcal{E}}}$ (3)
[…]
For a given atomic configuration, the BEC can be computed from any of the two expressions in Eq. (3). The (relative) efficiency of these approaches strongly depends on the system, on the chosen electronic-structure code, and on details of the implementations. Evidently, the expressions can be calculated by finite differences or by perturbation theory [32, 33]. While the latter is elegant and can be quite efficient especially for systems where symmetry can be exploited, the former ensures a natural handling of electronic-structure level of approximation, such as different density-functionals, relativistic corrections and other terms that are cumbersome to evaluate analytically. In any event, the calculation of $\bm{Z}^{*}$ always depends on the calculation of $\bm{\mu}$, and thus represents an added cost on the already costly electronic-structure calculation.
We aim to circumvent the expensive first-principles calculation of $\bm{Z}^{*}$ that enter Eq. (2) by developing a ML model that is trained only on $\bm{\mu}$ and that is autodifferentiable [34] with respect to its input quantities, the nuclear coordinates $\bm{R}$. Training on $\bm{\mu}$ presents some advantages, including the ease to produce data for this quantity with diverse electronic-structure methods and the guarantee that the acoustic sum rule [35] for $\bm{Z}^{*}$ in the model is preserved, making it strictly translationally invariant.
II.1 Dipoles in Aperiodic and Periodic Systems
While the calculation of the nuclear contribution $\bm{\mu}_{n}$ to $\bm{\mu}=\bm{\mu}_{e}+\bm{\mu}_{n}$ presents no issues, the electronic contribution $\bm{\mu}_{e}$ can be more cumbersome to calculate. In aperiodic systems, $\bm{\mu}_{e}$ can be obtained through a real-space integral involving the ground-state electronic density $n_{0}(\bm{r})$
[…]
In periodic, condensed-phase systems, $\bm{\mu}_{e}$ cannot be directly evaluated by the expression above, as restricting the integration domain to the primitive unit cell would yield boundary-sensitive results. The dipole is related to the polarization $\bm{P}$ of a periodic system, in that the polarization is the dipole per unit volume, $\bm{P}=\bm{\mu}/\Omega$. The modern theory of polarization (MTP) [36, 37, 38] provides a proper definition of this quantity in periodic systems, showing that the electronic contributions to $\bm{P}$ come from many-body effects associated to the geometric phase of the wavefunction. There are several different (but equivalent) ways to evaluate $\bm{\mu}_{e}$ within MTP, which are all related to a way of evaluating the Berry connection [37, 38, 39]. As a consequence, the components of $\bm{\mu}$ corresponding to each lattice vector $\bm{a}_{\alpha}$ can be defined only up to quantum of polarization $e\left|\bm{a}_{\alpha}\right|$ and the dipole along $\bm{a}_{\alpha}$ is
$\mu_{\alpha}\equiv\mu_{\alpha}\mod e|\bm{a}_{\alpha}|.$ (5)
Because of the congruence relation in Eq. (5), the dipole is better represented as a lattice of vectors [38], and only differences of $\mu_{\alpha}$ values can be well defined. If one imposes translational invariance on the dipole by enforcing $\bm{\mu}$ to assume the same value at translationally-equivalent periodic atomic sites, the function $\bm{\mu}(\bm{R})$ becomes discontinuous because this implies “switching branches” across the periodic boundaries, as shown in Fig. 1a. Instead, a smooth and continuous function is obtained by continuing to follow the same branch, (Fig. 1a). Since the dipole then assumes different values for translationally-equivalent atomic positions, $\bm{\mu}(\bm{R})$ does not have translation invariance with respect to the periodic boundary conditions.
Figure 1: Multi-valued dipoles and oxidation numbers. (a) The top panel shows a sketch of the DFT dipole values when folded on different branches: this leads to a discontinuous function that assumes exactly the same value for the same atomic environment ($x=0$ mod $|\bm{a}_{\alpha}|$). A single-valued ML approach can be successful if the training data and the predicted configuration lie in the “equilibrium” region (orange). The bottom panel shows the DFT dipole values lying on the same branch and leading to a continuous function. This requires a multi-valued ML model that is able to predict different values for the same atomic environment. This approach is also successful in the “far from equilibrium” region. (b) DFT values of the projection of the dipole along a given path for a water monomer (see Eq. (8)). The difference between the initial and final values (corresponding to equivalent configurations) correspond to the oxidation numbers of the atoms. (c) Analogous to b for LiNbO3.
In other words, a consequence of the multi-valued nature of $\bm{\mu}$ is that closed integrals do not vanish, but can take up (multiple) polarization quanta, depending on how often the periodic boundaries are crossed. We can quantify the change of the dipole $\Delta\bm{\mu}$ along a closed path $\gamma$ by a line integral. The result of this operation is non-trivial [40]:
$\Delta\bm{\mu}\left[\gamma\right]$ $=\,\oint_{\gamma}d\bm{\mu}=\oint_{\gamma}d\bm{r}\frac{\partial\bm{\mu}}{\partial\bm{R}}\biggr{\rvert}_{\bm{R}=\bm{r}}=e\bm{l},$ (6)
where $\bm{l}=\bm{a}\cdot\bm{n}$ is an element of the Bravais lattice of the system, $\bm{n}\in\mathbb{Z}^{3}$ and $\bm{a}$ is the matrix of lattice vectors $\bm{a}_{\alpha}$. This result shows that $\Delta\bm{\mu}=\bm{0}$ when no atom crosses the cell boundaries along $\gamma$. However, if $\gamma$ involves a displacement to a periodic replica at $\bm{R}+\bm{a}\cdot\bm{n}^{\prime}$ ($\bm{n}^{\prime}\in\mathbb{Z}^{3}$), the ratio (component-wise) between $\bm{n}$ and $\bm{n}^{\prime}$ has to be an integer $\mathcal{N}$ [40], i.e.
$\bm{n}=\mathcal{N}\bm{n}^{\prime}.$ (7)
This relation has been extensively discussed in by Jiang, Levchenko and Rappe in Ref. [40], where the authors related $\mathcal{N}$ to the atomic oxidation numbers. $\mathcal{N}$ does not depend on the path $\gamma$ as long as no metal-insulator transition occurs and it is defined for each atom of the system. Therefore, $\Delta\bm{\mu}$ can be non-zero if $\mathcal{N}\neq 0$. This relation also plays a fundamental role for ionic transport [41], where topological arguments relate transport to formal oxidation numbers [42].
Obtaining $\mathcal{N}$ for different atomic species from a DFT calculation is straightforward. One computes the dipole $\bm{\mu}$ along a path $\gamma$ where the $I^{th}$ atom is displaced from its position $\bm{R}_{I}$ to $\bm{R}_{I}+\bm{a}\cdot\bm{n}^{\prime}$ and evaluates $\mathcal{N}=\Delta\bm{\mu}\cdot\bm{l}/(e\left|\bm{l}\right|^{2})$, with $\bm{l}=\bm{a}\cdot\bm{n}^{\prime}$. In Fig. 1b and Fig. 1c we show this calculation graphically for a water molecule in a box and the ferroelectric solid LiNbO3, respectively. By defining
$\mathcal{N}(s)=\frac{\bm{\mu}(s)\cdot\bm{l}}{eN_{\rm at}|\bm{l}|^{2}}$ (8)
we can describe the more general case case where $N_{\rm at}$ atoms of the same oxidation state are displaced along the same direction. $\mathcal{N}_{I}$ is the difference between the values of $\mathcal{N}(s)$ upon crossing the periodic boundary. The values obtained for $\mathcal{N}_{I}$ are reported in Fig. 1b and c, and follow the expected oxidation numbers in these systems.
II.2 Machine-learning model for dipoles
The previous considerations lead to practical consequences when building a ML model for $\bm{\mu}(\bm{R})$ of a periodic system. Currently, equivariant neural networks (ENNs) yield the best accuracy to machine-learn properties of materials [43]. Most commonly, ENNs take as inputs the tuple $\left(Z;\bm{R},\bm{a}_{\alpha}\right)$, where $Z$ is the atomic number, and model the target quantity as a function of the interatomic distances $\Delta\bm{R}_{IJ}$, which are evaluated within the minimum image convention. We will call an atomic environment $\bm{\mathcal{A}}$ an element of the configuration space, which is isomorphic to a $3N$-dimensional torus $\mathbb{T}^{3N}$.
Going back to Fig. 1a, because such models require that the function to be learned is differentiable, the discontinuous curve which is obtained by matching $\bm{\mu}$ to assume the same value at equivalent $\bm{\mathcal{A}}$ cannot be learned in its full domain. If one is only interested in small atomic displacements close to equilibrium, a usual ML model will provide accurate results for $\bm{\mu}$ in this restricted space. However, as one moves into $\bm{\mathcal{A}}(\bm{R})$ that are far from equilibrium, the ML model will start to change its slope, in order to obey $\bm{\mu}(\bm{\mathcal{A}}(\bm{R}+\bm{l}))=\bm{\mu}(\bm{\mathcal{A}}(\bm{R}))$. In an autodifferentiable model, this results in $\bm{Z}^{*}$ and consequently forces that start to deviate from the correct value. If more data including environments far from equilibrium are added to the training set, the model will be unable to learn (see Supplementary Sections S2 for a practical example).
On the other hand, as one can also see in Fig. 1a, the multi-valued dipole, where $\bm{\mu}(\bm{\mathcal{A}}(\bm{R}+\bm{l}))\neq\bm{\mu}(\bm{\mathcal{A}}(\bm{R}))$, is a continuous smooth function. Learning this function, however, requires a modification to the usual architectures, to be able to describe this multi-valued nature of $\bm{\mu}$. In this work, we have slightly modified the MACE equivariant message-passing neural network [44, 45] to define a model $\tilde{\bm{\mu}}^{\text{MV}}$ which depends directly on $\bm{R}$ and $\bm{\mathcal{A}}$ as follows,
$\tilde{\bm{\mu}}^{\text{MV}}(\bm{R})=\tilde{\bm{\mu}}(\bm{\mathcal{A}})+e\sum_{I}\mathcal{N}_{I}\bm{R}_{I},$ (9)
where $I$ runs over all atoms of the system and $\tilde{\bm{\mu}}$ can be handled by the usual learning procedure, as it depends only on $\bm{\mathcal{A}}$. From the definition of $\tilde{\bm{\mu}}^{\text{MV}}$ and $\bm{Z}^{*}$ in Eqs. (3) and (9), we arrive at
$\bm{Z}^{*}_{I}=\frac{1}{e}\frac{\partial\tilde{\bm{\mu}}^{\text{MV}}}{\partial\bm{R}_{I}}=\frac{1}{e}\frac{\partial\tilde{\bm{\mu}}}{\partial\bm{R}_{I}}+\mathcal{N}_{I}\cdot\mathbbold{1}_{3\times 3},$ (10)
In the implementation we used in this work, we modified the output of the forward pass of MACE to include the second term on the right of Eq. (9). In this way, we always calculate $\tilde{\bm{\mu}}^{\text{MV}}$ in the training and use that also in the calculation of the loss-function. Training is successful and accurate even for datasets containing very out of equilibrium structures of H2O and LiNbO3, and $\bm{Z}^{*}$ obtained by autodifferentiation is also accurate (see Supplementary Section S2 and Section V).
The successes of previous ML models of dipoles and BEC are numerous [10, 14, 11, 46, 47, 18, 48, 16, 19, 9, 17]. They are successful because they either treat diffusive systems where the combined $\mathcal{N}$ (summing over all diffusing atoms) is zero, or address problems where the displacements along “ferroelectric” modes are not large and the system stays close to equilibrium along that coordinate. We note that models targeting to learn $\bm{Z}^{*}$ can also be problematic far from equilibrium if they miss the last term in Eq. (10) and they are not guaranteed to obey the acoustic sum rule for this quantity. The approach we present encompasses and extends these models to be generally applicable. This is for instance important for ionic diffusion [42] and strongly non-equilibrium phonon driving of ferroelectric materials, as we show below.
III Results
III.1 Dielectric Properties and Electrofreezing of Water
Figure 2: Liquid water properties from ML-based molecular dynamics simulations under applied electric fields. (a) Calculated real and imaginary parts of the dielectric permittivity at 300 K in the 1-135000 GHz frequency range (see text). The gray points correspond to experimental data at 20C from Ref. [21]. Dashed lines serve as guides to the eye. (b) IR absorption spectra of liquid water at 300 K and varying electric-field strengths. (c) Average of the polarization along the direction of applied electric fields of varying strengths obtained from simulations with classical nuclei (red) and quantum nuclei (blue). Dashed lines represent fits assuming a model of classical non-interacting dipoles. (d) OH bond length and O-O nearest neighbor distance with varying field strengths in simulations with classical (red) and quantum (blue) nuclei. Also shown is the self-diffusion coefficient of water calculated from the root-mean-square displacements of oxygen atoms (classical nuclei) at varying field strengths. Shaded areas in all panels correspond to the standard error.
We first apply the ML models we developed to study liquid water at room temperature and standard density. Water is of fundamental importance to life and, as such, it has been addressed numerous times in literature [49, 21, 50], with many peculiar features related to its dielectric properties and vibrational dynamics under applied electric fields still not fully resolved. The polarization of water results from a complex interplay between changes in the dipoles of individual water molecules and the correlated dynamics of multiple molecules in the liquid state. The results we present below were obtained by coupling the ML model we developed for the dipoles with a MLIP model for the energies and forces, see Section V.
Our implementation allows nanosecond long simulations and the application of dynamic and static electric fields, going well beyond the time-scales reached by existing ab initio techniques. It is thus possible to efficiently simulate the frequency-dependent dielectric function $\varepsilon(\nu)$ of liquid water, as shown in Fig. 2a, in the range of $0.001-135$ THz ($0.03-4500$ cm-1). Spanning this large frequency range was made even more efficient through the combination of two simulation techniques that are allowed by our methodology. Above 100 GHz, the real and imaginary parts of $\varepsilon(\nu)$ were calculated through the fluctuation-dissipation relation [51, 52, 10] (see Supplementary Section S3), while the region below 100 GHz was obtained by a more efficient non-equilibrium simulation technique [51] (see Section V). We could then simulate the onset of the Debye relaxation, which reduces the value of $\varepsilon(\nu)$ at frequencies larger than 10 GHz [21]. The results shown in Fig. 2a agree quite well with experimental data [21] shown in the same figure, despite a slight underestimation of the static dielectric constant, which we attribute to the training of the $\tilde{\bm{\mu}}^{\text{MV}}$ model in this case. As we describe in Section V, the training was effectively based on mid-sized water clusters. We expect that training on larger periodic structures will solve this issue, better reproducing previous ab initio results [53]. Besides these differences, the results substantiate that our approach is able to qualitatively capture the atomistic mechanisms of the Debye relaxation in water and should be a useful tool to settle the debate about the spatial extent of dynamical molecular correlation involved in this process, without having to resort to empirical potentials for large system sizes [50, 53].
In Fig. 2b we show the IR spectra of water under applied static electric fields, as obtained from the Fourier transform of dipole-autocorrelation time-series, simulated with molecular dynamics. The IR spectrum is closely related to ${\rm Im}\,\varepsilon(\nu)$ shown in Fig. 2a. The spectra in Fig. 2a are in excellent agreement with the ones recently reported in Ref. [17], which employed a similar ansatz of linear-coupling to the electric-field as we do here. With increasing field strength, there is a pronounced blue-shift of the libration band below $1000$ cm-1, a very small blue-shift of the bending band at $\approx 1700$ cm-1 and of the combination band at $\approx 2200$ cm-1, and a pronounced red-shift of the OH stretching band at $\approx 3600$ cm-1, effectively narrowing the spectral range. Such shifts are common features of the vibrational Stark effect in anharmonic systems [54]. The red-shift of the OH stretching band is due to a strengthening of H-bonds, evidenced by shorter O–O distances, as shown in Fig. 2d. The blue-shift of the libration band is due to the breaking of molecular rotational isotropy at increasing field strengths, as also discussed in Refs. [17, 23, 55]. Related to this blue-shift in the libration band, while the band we calculate is in excellent agreement with the one shown in Ref. [17] at all field strengths, both results do not contain the new peak just below $1000$ cm-1 at higher electric field strengths, reported by Cassone and coworkers and by Futera and English [23, 55], with ab initio MD simulations coupled to electric fields. This band is ascribed to a libration motion that is enhanced because of the preferential orientation of water along the electric field direction. As we also observe the preferential orientation of water molecules in our study (Fig. 2c), these differences must depend on the underlying methodology. We propose that the dynamics involved in the enhancement of this peak may be strongly dependent on the different underlying exchange-correlation functionals or on the inclusion of second-order electric-field coupling in the nuclear Hamiltonian, which are not explicitly considered in our work or in that of Ref. [17]. Because it is straightforward to augment our methodology to account for these terms, these considerations will be the subject of future work.
We now analyze the impact of nuclear-quantum-effects (NQE) together with the application of electric fields in water. To the best of our knowledge, the importance of these effects at higher electric field strengths has not been previously discussed. In Fig. 2c we report the average polarization of liquid water along the field direction and at varying field strengths $\mathcal{E}$, from classical-nuclei and path-integral molecular dynamics (see Section V). While at zero-field there is no net polarization, the interaction with the field causes a net polarization to appear along the field direction. The increase of the average polarization for $\mathcal{E}\gtrsim 0.04$ V/Å  is non-linear, in agreement with Ref. [17]. This behavior can be understood by analyzing the thermodynamics of an ensemble of classical non-interacting dipoles in an external field. The analytical expression of the partition function of this system can be found in textbooks [56], and the total polarization of the system is expressed as a Langevin function that depends on $\mathcal{E}$ and the molecular dipoles. By fitting to this model (dashed lines in Fig. 2c), we conclude that NQEs make water molecules easier to polarize upon electric field application. The saturation value of the average polarization is higher when including NQE, because the enthalpic term of interaction with the field can better compensate the thermal entropy of the liquid.
The impact of NQE on the average polarization correlates with the changes in structural properties of water, shown in Fig. 2d. With increasing $\mathcal{E}$ the OH bond length increases while the O-O nearest-neighbor distances decrease, consistent with the strengthening of the H-bonds. These effects are much more pronounced at $\mathcal{E}=0.15$ V/Å  when including NQE. The stronger electric fields thus shift the balance of the competing NQE in water [57], causing NQE to work towards considerably strengthening the H-bond interaction.
For the classical-nuclei simulations, we also calculated the water self-diffusion constant at different $\mathcal{E}$, shown in Fig. 2d. Consistent with the phenomenon of electrofreezing [23], we observe this coefficient to decrease from around 1.2 Å2/ps at zero-field to 0.3 Å2/ps at $\mathcal{E}=0.15$ V/Å. Based on the previous considerations, it is certain that NQE will also enhance this effect by causing a faster decrease in the diffusion coefficient with increasing field strength.
III.2 Phase Transition and Non-equilibrium Phonon Driving in LiNbO3
Figure 3: Phase transition and non-equilibrium properties of LiNbO3 from ML-based molecular dynamics simulations. (a) Thermal average of the dipole at varying temperature. The polarization shows a transition from a ferroelectric state (at low temperature) to a paraelectric state with a Curie temperature (red dashed line) close to the experimental value $T_{C}=1413$ K [26]. The polarization was computed as the dipole $\bm{\mu}$ per unit volume $\Omega$ along the direction $\left[111\right]$. (b) The polarization fluctuations $\delta\mu/\Omega$ shows a broadened peak around the Curie temperature as expected from a second-order phase transition simulated in a finite-system. Shaded areas correspond to the standard error over uncorrelated trajectories. (c) Ultrafast dynamics of the polarization of LiNbO3 under vibrational excitation with a monochromatic THz laser pulse with $\nu$ = 18 THz and FWHM = 188 fs and at varying maximum intensities. The $y$ axis shows the polarization with respect to the paraelectric state. The gray area represents the window of time during which the pulse is applied and the origin of time is at the center of the pulse.
Next, we show the generality of the ML approach developed here to address the ferroelectric to paraelectric phase-transition in solid LiNbO3, a widely employed ferroelectric perovskite with a Curie temperature of $T_{C}=1413$ K [25, 58, 26]. The study of this system presents a counterpart to the previous results on water, as polarization changes are predominantly induced by the large-amplitude motion of the Li+ ions. In this system, a single-valued model for $\bm{\mu}$ would readily fail. As we show in the Supplementary Section S2, when attempting to train a single-valued model on the dataset of LiNbO3 including structures where the Li+ atoms have been driven far from their equilibrium positions, the training is simply not successful.
We have evaluated the mean and the variance (fluctuations) of the dipole in the temperature range of $10$ to $1800$ K. The results are shown in Fig. 3a and b. Both plots show the typical behavior of a second-order phase transition [26]. The average value of the order parameter (the dipole) decreases to zero above the Curie temperature (Fig. 3a), while its fluctuations show a sharp peak at the same temperature (Fig. 3b). Finite size effects cause this peak to broaden with respect to the expected $\delta$-function. The analysis of the atomic displacements presented in Supplementary Section S4, which are in agreement with simulations reported in Ref. [59], confirm that the simulations indeed capture the ferroelectric to paraelectric phase-transition.
The phase-transition temperature inferred from Fig. 3a and b matches quite well the experimental Curie temperature $T_{C}$, which is likely to be partially fortuitous, because our simulations do not include the effects of thermal expansion. These effects are expected to increase $T_{C}$ by around 100 K [59]. Nevertheless, the reasons why the value we calculate is so close to the experimental result are that we properly account for all orders of anharmonic coupling between vibrational modes and that we accurately model its polarization changes, including its multi-valued nature. The absence of anharmonic couplings was previously shown to cause an underestimation of 250-300 K for $T_{C}$ [59, 60].
LiNbO3 has been at the center of active scientific debates due to the possibility of reversing its polarization using THz laser pulses which induce a non-thermal excitation of an IR-active mode $Q_{\rm IR}$ that strongly couples to a polarization-reversal mode $Q_{\rm P}$, as theoretically proposed in Ref. [61] and partially realized in Ref. [28]. To date, theoretical studies modeling this effect [62, 63, 29] were based on approximations to the PES which consider two or only a few phonon modes and the lowest-order anharmonic couplings between them. We have employed the ML models developed in this work to address this question, without relying on such approximations.
We have performed simulations where we applied a time-dependent monochromatic pulse with frequency $\nu=18$ THz enveloped with a Gaussian function with FWHM=188 fs and different maximum field intensities $\mathcal{E}_{\mathrm{max}}$ in the range $0.05-0.30$ V/Å. These values were chosen based on the reported experimental setups [28, 29] (see also Supplementary Section S5). The excitation is in close resonance with the 18.4 THz $Q_{\rm IR}$ mode, while the proposed $Q_{\rm P}$ mode lies at $7.4$ THz. These modes are depicted in Supplementary Section S6.
As shown in the Supplementary Sections S6 and S7, we observe the non-linear coupling of the $Q_{\rm IR}$ mode that is directly excited by the pulse at $\nu=18$ THz with the $Q_{\rm P}$ mode, which starts accumulating energy with a delay of some hundreds of fs. In addition we observe that only two other modes $\Gamma$-point modes, of A1 symmetry, at 8.2 and 10.0 THz get excited at later times. These modes are also IR active and, in particular the mode at 8.2 THz contributes to the polarization-reversal, attesting to the complex dynamical coupling in this system. In Fig. 3c we show the time-dependent variation of the polarization at times previous to, during (gray shaded area) and after the laser pulse is applied. The value of the polarization deviates from the thermal equilibrium value as soon the laser pulse starts acting on the system. We can recognize three different regimes, depending on the maximum intensity of the laser pulse. (i) At low intensities such as $0.05$ V/Å, the polarization is only slightly perturbed and the system remains ferroelectric. (ii) At higher intensities ($0.15$ and $0.20$ V/Å), the system is transiently driven to the paraelectric state (with zero polarization) during the laser pulse, but returns to the original ferroelectric state with a relaxation time of about $300$ fs, during which coherent dynamics can still be observed. (iii) At extremely high fields ($0.25$ and $0.30$ V/Å), the system transiently reaches higher values of reversed polarization but never fully switches, and after the pulse is switched off it transitions to the paraelectric state with incoherent dynamics because the amount of energy absorbed effectively heats the system above $T_{C}$.
We do not observe a full coherent switching of the polarization in LiNbO3. This is in agreement with experiments [28], which could not fully achieve a polarization reversal with this type of phonon driving. This has been attributed to the presence of depolarization fields [30], which prevent the full reversal, and which we can capture in the large unit cells used for the dynamics in this study. However, Fig. 3c shows that, within the pulse duration and at intense enough fields, it is possible to drive the system coherently to the paraelectric state. We propose that this is likely the explanation for the dip in the second-harmonic generation signal during the application of the field observed in the experiments reported in Ref. [28].
IV Conclusion
The results we have shown attest to the reliability and efficiency of the dipole ML model we developed. Together with a molecular dynamics protocol that captures the coupling of time-dependent and static electric fields with matter, it can deliver accurate properties and new physical insights for liquids and solids, both at equilibrium and driven far from equilibrium. This formulation thus extends the applicability realm of several ML models previously proposed, being applicable to isolated and periodic systems on the same footing and automatically satisfying charge conservation.
Regarding the specific applications we have shown, we would like to highlight a few points. In water, while the changes in the polarization behaviour due to NQE under applied electric fields are seemingly small, the magnitude of such changes can lead to vastly different state-points of phase transitions. We expect these effects to be particularly significant for superionic phases of water, which can be tuned by electric fields and by nanoconfinement [64, 65]. A study of the stability and dynamics of superionic phases under applied electric fields with ML methods would require a multi-valued dipole model such as the one developed here.
Indeed, we have shown that the necessity of employing a multi-valued ML model is paramount to capture phase transitions and general metastable and non-equilibrium states in LiNbO3. With an extension of our method to account for the coupling of electric fields with the lattice degrees of freedom, we also expect to be able to capture non-equilibrium dynamics of piezoelectric solids, considerably increasing the breadth of simulation tools in the area of vibrational ultrafast driving of materials.
In summary we expect that the simple technique we proposed, which incorporates the fundamental topological properties of the polarization of periodic systems into ML models, will bring significant advantages to the simulation of a range of liquids and solids that vastly surpass the examples shown here.
V Methods
The evolution of the nuclear equations of motion including time-dependent and independent applied electric fields was obtained by an implementation developed by us in the i-PI code, which was recently described in Ref. [66]. We refer the reader to that publication for technical details. All molecular-dynamics simulations were run with i-PI as the driving code communicating with the ML models, which provided all components of the force, and which are further described below.
parameters H2O LiNbO3
n. atoms 3 10
$a=b=c$ [Å] 4 5.48
$\alpha=\beta=\gamma$ [deg] 90 55.91
Table 1: Unit cell parameters used for the calculation of the oxidation numbers as shown in Section II.1.
species direction n. atoms $\mathcal{N}$
H2O
H $\left[0,\phantom{-}1,\phantom{-}0\right]$ 2 -1
O $\left[0,\phantom{-}1,\phantom{-}0\right]$ 1 -2
LiNbO3
Li $\left[1,-1,\phantom{-}0\right]$ 2 -1
Nb $\left[1,-1,\phantom{-}0\right]$ 2 -5
O $\left[1,\phantom{-}1,\phantom{-}1\right]$ 6 -2
Table 2: Computational details used to evaluate the oxidation numbers (as presented in Section II.1) of all the atomic species in water and LiNbO3. The table reports the direction, on the basis of the lattice vectors, along which the atoms were displaced, the total number of atoms of that species, and the evaluated oxidation number.
quantity unit H2O LiNbO3
energy meV/atom 1.10* 0.85
forces meV/Å 31.19* 27.51
dipole mD/atom 3.12* 11.38
Table 3: Performance of the trained models for liquid water and LiNbO3. The values marked with an asterisk represent RMSE on the training dataset, all other values represent RMSE on the test set. See Supplementary Section S8 for further details and information.
V.1 Water
We have evaluated the oxidation numbers of the H and O atoms of a single water molecule following Section II.1. A cubic simulation box of $4$  (whose parameters are reported in Table 1) was used. Figure 1b was obtained considering $100$ configurations, where the H atoms were uniformly displaced along the direction reported in Table 2 from the relaxed initial configuration, while the O atom was fixed, and vice-versa. The DFT calculations were performed with the revPBE functional, no spin polarization, a $3\!\times\!3\!\times\!$ 3 × 3 $\bm{k}$-grid, atomic ZORA relativistic treatment [67], and the intermediate basis set in the FHI-aims [68] code. The polarization components were calculated [69] on a denser $\bm{k}$-grid, namely $40\!\times\!10\!\times\!10$ for the component along $\bm{a}_{1}$, $10\!\times\!40\!\times\!10$ for $\bm{a}_{2}$, and $10\!\times\!10\!\times\!40$ for $\bm{a}_{3}$.
The machine-learning interatomic potential (MLIP) for liquid water was trained with a MACE [44, 45] model on the data provided in Ref [70], i.e. energy and forces computed using the revPBE0 functional with Grimme D3 [71, 72] dispersion corrections. We have trained a MACE MLIP with a cutoff radius of $6$, $2$ layers, $64$ embedding channels, $l=3$ as the highest spherical harmonics, a maximum of $l=0$ for each message, and correlation order $3$ at each interaction layer.
The dataset for training the dipole model was obtained from 3 NVT simulations (stochastic-velocity-rescaling thermostat SVR  [73]; $\tau=100$ fs; $T=330,360,390$ K) with i-PI [66] and the trained MLIP as a force provider. We ran $100$ ps for each run with a time-step of $1$ fs. The simulation box was cubic with lattice vectors of length $14.9365$  and $111$ molecules, corresponding to a density of $0.9965$ g/m-3. The farthest point sampling (FPS) algorithm, as implemented in the librascal library [74] and based on SOAP descriptors as reported in Ref. [75] was used to select the $1000$ most diverse structures from all trajectories.
We used the structures in the unit-cell in an aperiodic setting (effectively water droplets) and computed the dipole using Eq. (4) with the revPBE functional. This was done solely to increase efficiency of data generation. We have split the data with a ratio of $80/20$ for training and validation, after having separated 20% of the full dataset for testing. The dipole model was trained using a modified MACE architecture developed by us and available in Ref. [76], with a cutoff radius of $6$, $2$ layers, $14\!\times\!0e\!+\!4\!\times\!$ 1o as irreps for the hidden node states, $l=2$ as the highest spherical harmonic, correlation order of $3$ at each layer. We added the oxidation numbers contribution according to Eq. (9) in the training. The performance of these models are reported in Table 3.
The IR absorption spectra of liquid water and field-dependent self-diffusion coefficients (Fig. 2a and d) were obtained by first equilibrating the system for $100$ ps at $300$ K (Langevin thermostat, $\tau=50$ fs, $dt=0.25$ fs) and then applying the constant electric fields following Eqs. (1) and 2. The MLIP together with the ML dipole model were used to obtain the forces. We picked $16$ random snapshots from this trajectory and used them to start 16 NVE trajectories of $40$ ps each. The IR absorption spectra were obtained from the Fourier transform of the autocorrelation of the dipole derivatives [77]. The diffusion coefficient was obtained by computing the mean square displacement (MSD) of the O atoms for each time-step and taking the average value of its time-derivative on the last $10$ ps, when the behavior is already linear with time.
The high-frequency part of the dielectric susceptibility was computed with the fluctuation-dissipation relation of the dipole and dipole-derivative equilibrium correlation function [51, 52, 10], using the same data as for the infrared absorption spectrum with zero electric field. The low-frequency part of the dielectric susceptibility was computed using the non-equilibrium direct electric field method described in Ref. [51]. We simulated NVT trajectories of $1-2$ ns (SVR thermostat, $\tau=100$ fs, $dt=0.5$ fs) including a time-dependent monochromatic external electric field of $\mathcal{E}_{\rm max}=0.04$ V/Å, with frequencies of $1,3,10,30,100$ and $200$ GHz. The time-dependent behavior of the (induced) polarization at each frequency was fitted with a function $P_{\rm ind}\cos\left(2\pi\nu-\phi\right)$ to obtain the parameters $P_{\rm ind}$ and $\phi$. The introduction of the 3rd and 5th harmonics in the fit, as done in Ref. [51], resulted in negligible corrections to the fitted values. The real and imaginary parts of the dielectric susceptibility were then computed as $P_{\rm ind}\cos\left(\phi\right)/\varepsilon_{0}\mathcal{E}_{\rm max}$ and $P_{\rm ind}\sin\left(\phi\right)/\varepsilon_{0}\mathcal{E}_{\rm max}$ respectively.
The dipole and structural data in Figs. 2c and d were derived from $16$ NVT trajectories for each different electric field of $10$ ps each. The static structural properties related to electrofreezing were also calculated including nuclear quantum effects with path integral molecular dynamics (PIMD). We ran the same number and length of trajectories for each electric field intensity, for 30 ps, using $8$ ring-polymer replicas, a time-step of 0.25 fs, and the PIGLET thermostat [78].
V.2 LiNbO3
The oxidation numbers of the atoms in LiNbO3 presented in Fig. 1c were computed by using the fully relaxed unit cell whose parameters are reported in Table 1 with space group $R3c$. The atoms were displaced along the directions reported in Table 2. We employed the PBEsol functional, a $\bm{k}$-grid of $3\times 3\times 3$, no spin polarization, atomic ZORA relativistic treatment, and the intermediate basis set in FHI-aims. All atoms of the same species were displaced together.
The LiNbO3 dataset for training the MACE MLIP and dipole models was obtained from $3$ ab initio NVT trajectories of 20 ps at 500 K, 1000 K, 1500 K (SVR thermostat, $\tau=100$ fs, $dt=2$ fs) with a simulation box containing $30$ atoms and lattice vectors $a,b,c=5.14,5.14,13.82$  and angles $\alpha,\beta,\gamma=90^{\circ},90^{\circ},120^{\circ}$. We used the FPS algorithm to select $5000$ structures. For these structures, we computed the polarization using the Berry-phase formalism implemented in FHI-aims [69]. The values of the dipoles have been branch-matched as described in Supplementary Section S9.
The MLIP consists of a MACE model with a cutoff radius of $6$, $2$ layers, $l=2$ as the highest spherical harmonics, a maximum of $l=0$ for each message, correlation order $3$ at each layer. The parameters for the MACE dipole model were the same as the ones used for water, but adjusting for the oxidation numbers of LiNbO3. The final errors on energies, forces and dipoles for these models are reported in Table 3.
The simulations regarding the ferroelectric to paraelectric phase transition in Section III.2 were obtained by using a $4\!\times\!4\!\times\!$ 4 × 4 supercell with $640$ atoms, lattice vectors of length $21.94$  and angles between them of $55.83^{\circ}$. For each temperature we simulated $10$ independent NVT runs (Langevin thermostat, $\tau=100$ fs, $dt=1$ fs). We used the trained dipole model to evaluate the dipole for each step.
The simulation with a time-dependent laser pulse, whose results are reported in Fig. 3c, were run with the same $4\!\times\!4\!\times\!$ 4 × 4 supercell as used above, and were run in the NVE ensemble, starting from a thermalized trajectory. We used a time-step of 0.1fs to ensure stability of the integrator with a time-dependent Hamiltonian. For each pulse considered, we ran 16 independent trajectories starting from randomly sampled geometries equilibrated at 300 K. The average and standard deviation at varying time shown in Fig. 3c correspond to the statistical analysis over these independent trajectories.
Acknowledgments
E.S. thanks Federico Ernesto Mocchetti for the mathematical background regarding the topological aspects of the multi-valued ML models and Shubham Sharma for discussions and help with training MACE models. We thank George Trenins for a critical read of the paper draft and insightful discussions. We thank Vasily Artemov for sending us a table with the experimental datapoints of the dielectric permittivity of water. We also thank Michael Fechner for insightful discussions about phonon dynamics in LiNbO3.
Author Contributions
M.R. and E.S. conceptualized and designed the project. E.S., C.C. and M.R. discussed and refined the theory. E.S. implemented the models and conducted the calculations, with supervision from M.R.. E.S. and M.R. analysed the results. M.R., E.S. and C.C. wrote the manuscript.